# references.bib {refsection}

### v3

@@
Abstract
Importance: Artificial Intelligence (AI) for health has the potential to significantly change and improve healthcare. However in most African countries, identifying culturally and contextually attuned approaches for deploying these solutions is not well understood.
Objective: To bridge this gap, we conduct a qualitative study to investigate the best practices, fairness indicators, and potential biases to mitigate when deploying AI for health in African countries, as well as explore opportunities where artificial intelligence could make a positive impact in health.
Design, Setting, and Participants: We used a mixed methods approach combining in-depth interviews (IDIs) and surveys. We conduct 1.5-2 hour long IDIs with 50 experts in health, policy, and AI across 17 countries, and, through an inductive approach, we conduct a qualitative thematic analysis on expert IDI responses. We administer a blinded 30-minute survey with case studies to 672 general population participants across 5 countries in Africa (Ghana, South Africa, Rwanda, Kenya and Nigeria), and analyze responses on quantitative scales, statistically comparing responses by country, age, gender, and level of familiarity with AI. We thematically summarize open-ended responses from surveys.
Main Outcome(s) and Measure(s): Our primary outcomes are to understand and compare general population and expert perceptions across the following axes: (1) current health inequities and general thoughts on addressing inequities; (2) perceptions around AI and healthcare and opportunities for development; (3) perceived effects of colonial history on AI; (4) fairness and bias considerations for machine learning in African contexts; and (5) community-driven solutioning for representative data generation and design and deployment of models.
Results: Our results find generally positive attitudes, high levels of trust, accompanied by moderate levels of concern among general population participants for AI usage for health in Africa. This contrasts with expert responses, where major themes revolved around trust/mistrust, ethical concerns, and systemic barriers to integration, among others. This work presents the first-of-its-kind qualitative research study of the potential of AI for health in Africa from an algorithmic fairness angle, with perspectives from both experts and the general population.
Conclusions and Relevance: We hope that this work guides policymakers and drives home the need for further research and the inclusion of general population perspectives in decision-making around AI usage.
1 Introduction
Artificial Intelligence (AI) has demonstrated significant contributions into various aspects of human lives. One of those areas of AI, and in particular data-driven Machine Learning (ML) solutions, is in health. While there is the potential for healthcare to be revolutionized by AI, overlooking geo-cultural and contextual knowledge could disproportionately and negatively impact human lives. Given that most of the historical data, models and problems are explored in Global North context, Global South geo-contexts are relatively underexplored. This study aims to help fill this gap by investigating the best practices, fairness indicators, and potential biases to mitigate when deploying AI for health in African countries, as well as explore opportunities where artificial intelligence could make a positive impact in health.
@@ 1 Introduction
Increasingly, advancements in AI technologies are having major impacts across a variety of healthcare delivery and services, such as improved diagnostics and precision medicine, predictive analytics and preventative care, personalized treatment, telemedicine, remote monitoring, and improved healthcare accessibility [Topol2019, Goldberg2024, patel2017correlating, Azizi2021, Singhal2023, jiang2017artificial, Yagnik2024]. More research is actively being done to discover the best use cases for AI in healthcare along with how to actualize those use cases. However, much of this work focuses on high income countries (HICs), and relies on assumptions about healthcare systems, practices, and health attitudes, which in many instances may be understood as belonging to distinctly Global North contexts [Wahl2018]. Conversely, in the context of low-and-middle income countries (LMICs), it is often the case that healthcare systems, contexts, practices, and prevailing health attitudes are importantly different from HICs in the Global North [Yu2024].
@@
Figure 1: Nteasee study overview of participant distribution within Africa, methods and summary of findings. Experts residing in countries outside of Africa were also interviewed as was one expert in Mauritius, though this is not reflected in the figure.
Related work:
To address this issue, a slew of recent research has sought to advance our understanding of the challenges and solutions for successful development and integration of AI into Global South health systems. At the cross-regional level, (Okolo2022; Kong2023; Chakraborty2024; Zuhair2024) advance insights from Global South contexts for improving the AI health regulatory landscape and for integrating AI into underserved and under-resourced health systems. At the regional level Sambasivan2021 have observed that mainstream conceptions of algorithmic fairness for model evaluation (both in health contexts and beyond) require retailoring when considering AI solutions in India. Bhatt et al. Bhatt2022 charted a framework for re-contextualizing natural language processing (NLP) fairness research in an Indian context, with the caveat that this can apply to other Global South geo-contexts. Concerning the African region specifically, research for successful AI integration into health systems traverses several subject-matter areas. The question of AI policy readiness and development Arakpogun2021; Gwagwa2021; otaigbe2022scaling; Etori2023; Diallo2024 have seen much attention recently, with some common themes being the need for robust data infrastructure, the importance of context-specific AI readiness assessments, and the critical role of international collaboration and capacity building in ensuring equitable AI development across the African continent. Better understanding challenges and solutions for AI in African health systems has also been a major point of focus, cutting across a range of topics such as addressing health inequities Qoseem2024, enhancing existing healthcare resources and infrastructure Owoyemi2020; Oladipo2024; Tshimula2024, and improving healthcare accessibility Sallstrom2019; Wakunuma2022; Mbunge2023, especially for under-resourced rural areas Guo2018; Arefin2024. In a related vein, there is a growing interest in the prospect of algorithmic colonization, where imported AI technologies from the Global North fail to align with local realities and even contribute to the perpetuation of Global South disparities Mohamed2020; Birhane2020; Asiedu2024. The concern over algorithmic colonization highlights the growing importance of developing decolonial approaches to AI integration Adams2021; Ayana2024, which are not only technically sound but also optimized for cultural attunement and equitable access.
The current literature underscores both the potential and challenges of AI in transforming healthcare in the African region. Expert stakeholders have previously been solicited to help identify conditions for developing AI solutions tailored to African contexts Lopes2021. Nevertheless, there remains a critical need to further include and grow the range of experts involved in building our understanding of policy and development considerations for successful AI integration in African health systems. Similarly, it is vital to also include the perspectives of general population stakeholders, who represent the ones ultimately directly affected by the use of AI health technologies.
Our contributions:
@@
In this study, we set out to explore what considerations may be relevant to deploying AI health solutions in low-and-middle income countries (LMICs) in Global South contexts. In the vein of exploring considerations for the deployment of AI health solutions in the Global South, our specific and original research focus is on the deployment of AI health solutions in African countries. The research questions we seek to answer include (1) what are currently the most significant determinants of health inequities, and perspectives on addressing them?; (2) what are present perceptions around AI and healthcare and opportunities for development?; (3) what are the perceived or envisioned effects of colonial history on AI?; (4) what are the fairness and ML bias considerations for African contexts specifically?; and (5) what are community-driven approaches to representative data generation and design and deployment of models?
@@
In order to answer these questions, we interviewed experts across the machine learning, healthcare, and health equity spheres, working within or around AI and health in African countries. In doing so, we drew out their insights across a variety of themes and dimensions involved in developing, deploying, and regulating AI solutions for health in African countries. We also deployed survey questionnaires with general population participants from five countries in Africa–Ghana, Rwanda, Nigeria, Kenya and South Africa–interrogating their attitudes towards AI health solutions. We compiled and analyzed these insights through an inductive approach, presenting them for a multi-stakeholder audience, namely technology and health policymakers, healthcare professionals, and machine learning developers and deployers.
@@
This work advances a picture of the landscape of AI for health in African countries, one which we hope may aid stakeholders in navigating the path towards efficient and impactful integration of AI health solutions across the African region. This being said, we acknowledge the limitations of our study and the insights it offers, among them the circumscribed range of African nations and linguistic and cultural contexts we investigate. Our aim through this study is to advance the discussion on the integration of AI solutions within such contexts, and beyond this to galvanize further critical research supporting the successful integration of AI health solutions across the Global South, and in doing so advance the discourse on AI, equity, and global health.
2 Methods
@@
Table 1: Participant Demographics for Experts IDI and General Population Survey Respondents
@@
We conducted a survey of general population (Gen Pop.) participants and in-depth interviews (IDIs) with pre-interview surveys of expert participants to understand perceptions of AI for health in Africa. Further details of the two arms of the study are in the subsequent sections. We analyze the survey and interview results independently and also analyze the intersections of the outcomes, comparing general population viewpoints to expert viewpoints.
2.1 General population surveys
2.1.1 Survey overview and participants
@@
2.1.2 Survey procedure
Respondents were first screened for eligibility. We asked eligible respondents to provide demographic information and then proceeded to ask questions about understanding of and familiarity with AI and ML and its benefits, trust in AI interventions for health, algorithmic fairness and biases, and intersections between AI, health, and colonialism. Additionally, we conducted case studies - providing hypothetical scenarios of a user interacting with an AI chatbot for health to understand conditions for acceptance, useability, and believability. We provided versions of these scenarios that were augmented to include location, gender and religion, to understand whether the above metrics changed with additional contextual responses. Finally we asked respondents to indicate specific non-personal hypothetical questions they would ask an AI chatbot about health. The full set of questions asked in the survey can be found in Appendix A.1.1.
2.1.3 Statistical analysis
@@
2.2 Expert in-depth interviews
2.2.1 Procedure
@@
We reached out to people in our networks using an email draft which included a link to fill out a participation interest survey for the interviews. We also shared the study description and fliers with a link/QR code to the form in AI communities and group chats, and on social media pages including LinkedIn. Participants were recruited between October 24 2023 and January 30th 2024.
@@
Participant eligibility criteria: As this part of the study specifically focused on experts, we recruited participants who had expertise in AI and health, practicing health providers, public health researchers, health policy experts, and entrepreneurs. We had 445 interested participants sign up to take part in the study. We went through five rounds of short listing to narrow down the participants based on responses in the participant interest survey. Participants were selected based on (1) years of experience, prioritizing >10 years of experience, followed by 5+ years of experience in their respective fields, (2) field of practice–we ensured there was a balanced representation of clinicians, public health, and AI researchers, (3) focus of research/work on Africa, and (4) country of residence, prioritizing African residents. We also included a few non-African residents who were African diaspora and/or had >10 years of experience working in Africa. Demographics of the participants as well as relevant criteria such as years of experience can be found in Table 1. Note that expertise in AI Trust/safety/responsible AI was not a selection criteria for experts, and none of the experts we interviewed indicated specifically that they were in the field of AI Trust/safety/responsible AI.
@@
Once selected to participate, we reached out to participants to confirm participation, and they were provided the appropriate informed consent for their geographic location. After agreeing to informed consent, the participants filled out a brief demographic survey. The interview was then scheduled and then conducted online over Google Video Conference (GVC) and audio-recorded. There were two individual moderators throughout the total qualitative interviews, and each interview was conducted with 1 moderator and 1 participant. During the interview, a moderator engaged with the participants in in-depth, semi-structured interviews that lasted approximately 60-90 minutes. Before each interview, the interviewer thoroughly explained the interview process. They emphasized the confidentiality of the participants’ responses and reminded them that the interview was being audio-recorded, and addressed any questions or concerns raised by the participants. After each interview, the voice recordings were saved as MP3 files and no identifying data (i.e., participants’ imagery) was retained. Voice recordings were transcribed by a vendor (D’ Well), and the transcripts stored in a secure data bucket without any PII. The full interview guide can be found in Appendix A.1.2.
We chose a semi-structured interview format, as opposed to a focus group approach, because interviews enable deeper exploration of individual perspectives and experiences that help mitigate the risk of conformity or social desirability bias that often occur in focus groups Creswell2017. This approach fosters a more accurate understanding of the subject matter, particularly around topics that explore complex and potentially controversial issues like AI in healthcare Kruger2018. Additionally, semi-structured interviews offer greater flexibility in scheduling and can be tailored to the specific expertise and interests of each healthcare worker or ML expert.
@@
The focus of the interviews explored the participants’ perspectives on (a) equity in healthcare; (b) historical biases in health (HBH); (c) machine learning in healthcare (MLH); (d) machine learning fairness framing and applications (MLF); (e) stereotypes, sensitive attributes, and vulnerable sub-populations (SV); and (f) an understanding of existing definitions of fairness (FD). Further, we explored (g) perceptions of colonialism and the Global North/South divide (CGNS). Semi-structured interviews have pre-determined questions that allow for flexibility in the interviewee’s responses Schensul1999. This approach allowed each of the moderators to probe and explore topics that emerged during the course of the interview with each participant.
2.2.2 Reflexitivity
Reflexivity is an important element of ensuring trustworthiness in qualitative research, particularly when the research team comprises individuals from diverse backgrounds and experiences which inevitably shape their interpretations of the data Anzul1991; Fonow1991. Therefore, it was crucial for the researchers involved in the qualitative analysis phase to critically reflect on their own biases and assumptions throughout the research process. Our diverse backgrounds and experiences influenced how we interpreted the data, bringing unique perspectives and insights to the analysis.
This ongoing self-awareness enhanced the credibility of the findings by ensuring that interpretations were grounded in the data rather than in the researchers’ preconceived notions. Moreover, by openly acknowledging our positionality, the researchers invited scrutiny and dialogue amongst one another to further strengthen the trustworthiness of the analysis during team meetings and data discussions. The process of reflexivity fosters transparency, acknowledges subjectivity, and ultimately enhances the rigor and credibility of qualitative research. To this end, the team met either weekly or bi-weekly to discuss the data process, the insights that were emerging, and engaged in critical discussion to ensure that the analysis codes accurately reflected the data. Two researchers work external to Google, but have a combined extensive background in bioethics, philosophy, and AI for social good in Africa. Five of the other researchers work internally at Google and work across the company in various departments. One has an academic background in Sociology and Cultural Studies, one has a background in biomedical engineering, AI and global health, three have backgrounds in machine learning, statistics and algorithmic fairness in machine learning. We also varied in age, race, national origin, and sexual orientation which influenced our experiences with, and perceptions of, health and AI. Two of the co-first authors are of African origin, as is one of the additional authors. Three of the study’s authors are of South Asian and/or Middle Eastern origin. Three of the study’s authors are of North American origin.
During our analysis and interpretation of the data, “self-situating” was crucial for our analytical process, even though it was neither feasible nor desirable to explicitly include our personal experiences in the analysis process (Frow2000). Guest2012 found that collaboration among individuals with diverse experiences enriches analysis of the interview data by offering multiple perspectives and backgrounds.
2.2.3 Data Analysis

### v4

# Nteasee: Understanding Needs in AI for Health in Africa - A Mixed-Methods Study of Expert and General Population Perspectives

@@
Abstract.
Artificial Intelligence (AI) for health has the potential to significantly change and improve healthcare. However in most African countries, identifying culturally and contextually attuned approaches for deploying these solutions is not well understood. To bridge this gap, we conduct a qualitative study to investigate the best practices, fairness indicators, and potential biases to mitigate when deploying AI for health in African countries. Additionally, we explore opportunities where AI could make a positive impact in health. We use a mixed-methods approach combining in-depth interviews (IDIs) and surveys. We conduct 1.5-2 hour long IDIs with 50 experts in health, policy, and AI across 17 countries, and, through an inductive approach, we conduct a qualitative thematic analysis on expert IDI responses. We administer a blinded 30-minute survey with case studies to 672 general population participants across 5 countries in Africa (Ghana, South Africa, Rwanda, Kenya and Nigeria), and analyze responses on quantitative scales, statistically comparing responses by country, age, gender, and level of familiarity with AI. We thematically summarize open-ended responses from surveys. Our results find generally positive attitudes, high levels of trust, accompanied by moderate levels of concern among general population participants for AI usage for health in Africa. This contrasts with expert responses, where major themes revolved around trust/mistrust, ethical concerns, and systemic barriers to integration, among others. This work presents the first-of-its-kind qualitative research study of the potential of AI for health in Africa from an algorithmic fairness angle, with perspectives from both experts and the general population. We hope that this work guides policymakers and drives home the need for further research and the inclusion of general population perspectives in decision-making around AI usage.
algorithmic fairness, bias, health equity, policy, need-finding, human experiences, norms, practices, participatory design
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732160isbn: 979-8-4007-1482-5/2025/06ccs: Computing methodologies Artificial intelligenceccs: Social and professional topics Computing / technology policyccs: Human-centered computing
1. Introduction
@@
Increasingly, advancements in AI technologies including large language models (LLMs) are promising to have significant impact across a variety of healthcare delivery and services, such as improved diagnostics and precision medicine, predictive analytics and preventative care, personalized treatment, telemedicine, remote monitoring, and improved healthcare accessibility (Topol, 2019; Goldberg et al., 2024; Patel et al., 2017; Azizi et al., 2021; Singhal et al., 2023; Jiang et al., 2017; Yagnik et al., 2024). More research is actively being done to discover the best use cases for AI in healthcare along with how to actualize those use cases through research examining human experiences, needs, perceptions and interactions. However, much of this work focuses on high income countries (HICs), and relies on assumptions about healthcare systems, practices, and health attitudes, which in many instances may be understood as belonging to distinctly Global North contexts (Wahl et al., 2018). Conversely, in the context of low-and-middle income countries (LMICs), it is often the case that healthcare systems, contexts, practices, and prevailing health attitudes are importantly different from HICs in the Global North (Yu and Zhai, 2024). While there is the potential for AI to revolutionize healthcare, overlooking geo-cultural and contextual knowledge could disproportionately and negatively impact human lives. This work aims to help fill this gap by investigating the best practices, fairness indicators, and potential biases to mitigate when deploying AI for health in African countries, as well as explore opportunities where AI could make a positive impact in health.
@@
Figure 1. Nteasee* study overview of participant distribution within Africa, methods and summary of findings. Experts residing in countries outside of Africa were also interviewed as was one expert in Mauritius, though this is not reflected in the figure.
* Nteasee is an Akan Twi word from Ghana which means “to understand”. It is also an Adinkra symbol (visual illustrations that represent proverbs, concepts and aphorisms) which stands for cooperation and understanding (Adjei, 2018).
Contributions:
@@
In this study, we set out to explore what considerations may be relevant to deploying AI health solutions in low-and-middle income countries (LMICs) in Global South contexts. In that vein, our specific and original research focus is on the deployment of AI health solutions in African countries. The research questions we seek to answer include (1) what are currently the most significant determinants of health inequities, and perspectives on addressing them?; (2) what are present perceptions around AI and healthcare and opportunities for development?; (3) what are the perceived or envisioned effects of colonial history on AI?; (4) what are algorithmic fairness and bias considerations for African contexts specifically?; and (5) what are community-driven approaches to representative data generation and design and deployment of AI systems?
@@
In order to answer these questions, we interviewed experts across the machine learning, healthcare, and health equity spheres, working within or around AI and health in African countries. In doing so, we drew out their insights across a variety of themes and dimensions involved in developing, deploying, and regulating AI solutions for health in African countries. We also deployed survey questionnaires with general population participants from five countries in Africa–Ghana, Rwanda, Nigeria, Kenya and South Africa–interrogating their attitudes towards AI health solutions including the use of case studies around large language models. We compiled and analyzed these insights through an inductive approach, presenting them for a multi-stakeholder audience, namely technology and health policymakers, healthcare professionals, and machine learning developers and deployers. To our knowledge, this is the first study that utilizes rigorous mixed-methods research with multinational, multidisciplinary experts and general population participants on AI for health in the Global South and to provide specific, actionable, best practices, and geo-contextual fairness and bias considerations.
Our work offers several novel insights regarding specific contextual challenges and opportunities for AI use in health in Africa. We suggest methods to build trust in AI systems to support adoption in the African context. Additionally, we identify that while experts have concerns about AI which may be indeed valid, general population participants are more accepting of AI adoption, emphasizing the need for policy makers to consider general population perspectives. This finding challenges the common assumption that the general population agrees with expert notions of data colonialism, fairness, and use of internationally developed AI tools.
This work advances a picture of the landscape of AI for health in African countries, delineating experience, needs, and perceptions of AI tools, which we hope may aid stakeholders in navigating the path towards efficient and impactful evaluation, governance and integration of AI health solutions across the African region.
@@
This being said, we acknowledge the limitations of our study and the insights it offers, among them the circumscribed range of African nations and linguistic and cultural contexts we investigate. Our aim through this study is to advance the discussion on the integration of AI solutions within such contexts, and beyond this to galvanize further critical research supporting the successful integration of AI health solutions across the Global South, and in doing so advance the discourse on AI, equity, and global health.
2. Related Work
Given that most of the historical data, models and problems in AI for healthcare are explored in Global North contexts, Global South geo-contexts are relatively underexplored. To address this issue, a slew of recent research has sought to advance our understanding of the challenges and solutions for successful development and integration of AI into Global South health systems. At the cross-regional level, Okolo (2022); Kong et al. (2023); Chakraborty and Karhade (2024); Zuhair et al. (2024) advance insights from Global South contexts for improving the AI health regulatory landscape and for integrating AI into underserved and under-resourced health systems. At the regional level, Sambasivan et al. (2021) have observed that mainstream conceptions of algorithmic fairness for model evaluation (both in health contexts and beyond) require retailoring when considering AI solutions in India. Bhatt et al. (Bhatt et al., 2022) charted a framework for re-contextualizing natural language processing (NLP) fairness research in an Indian context, with the caveat that this can apply to other Global South geo-contexts. Concerning the African region specifically, research for successful AI integration into health systems traverses several subject-matter areas. The question of AI policy readiness and development (Arakpogun et al., 2021; Gwagwa et al., 2021; Otaigbe, 2022; Etori et al., 2023; Diallo et al., 2024) have seen much attention recently, with some common themes being the need for robust data infrastructure, the importance of context-specific AI readiness assessments, and the critical role of international collaboration and capacity building in ensuring equitable AI development across the African continent. Better understanding challenges and solutions for AI in African health systems has also been a major point of focus, cutting across a range of topics such as addressing health inequities (Qoseem et al., 2024), enhancing existing healthcare resources and infrastructure (Owoyemi et al., 2020; Oladipo et al., 2024; Tshimula et al., 2024), and improving healthcare accessibility (Sallstrom et al., 2019; Wakunuma et al., 2022; Mbunge and Batani, 2023), especially for under-resourced rural areas (Guo and Li, 2018; Arefin, 2024). In a related vein, there is a growing interest in the prospect of algorithmic colonization, where imported AI technologies from the Global North fail to align with local realities and even contribute to the perpetuation of Global South disparities (Mohamed et al., 2020; Birhane, 2020; Asiedu et al., 2024). The concern over algorithmic colonization highlights the growing importance of developing decolonial approaches to AI integration (Adams, 2021; Ayana et al., 2024), which are not only technically sound but also optimized for cultural attunement and equitable access.
The current literature underscores both the potential and challenges of AI in transforming healthcare in the African region. Previous work has solicited expert stakeholders to help identify conditions for developing AI solutions tailored to African contexts (Lopes et al., 2021). Nevertheless, there remains a critical need to further include and grow the range of experts involved in building our understanding of policy and development considerations for successful AI integration in African health systems. Similarly, it is vital to also include the perspectives of general population stakeholders, who represent the ones ultimately directly affected by the use of AI health technologies.
3. Methods
@@
Table 1. Participant Demographics for Experts IDI and General Population Survey Respondents
@@
We conducted a survey of general population (Gen Pop.) participants and in-depth interviews (IDIs) with pre-interview surveys of expert participants to understand perceptions of AI for health in Africa. While health is a broad area and it is important to perform this study across different specific sub-specialties such as radiology, oncology, etc, we focus on a broad approach of health due to the paucity of original mixed-methods research in this space. This approach allowed the expert interviewees the flexibility to discuss which specific applications of AI they believe are most urgent and needed in the African context. Further details of the two arms of the study are in the subsequent sections. We analyze the survey and interview results independently and also analyze the intersections of the outcomes, comparing general population viewpoints to expert viewpoints.
3.1. General population surveys
3.1.1. Survey overview and participants
@@
3.1.2. Survey procedure
Respondents were first screened for eligibility. We asked eligible respondents to provide demographic information and then proceeded to ask questions about understanding of and familiarity with AI and ML and its benefits, trust in AI interventions for health, algorithmic fairness and biases, and intersections between AI, health, and colonialism. Additionally, we studied use cases of large language models (LLMs) due to their emerging relevance. Towards this, we conducted case studies - providing hypothetical scenarios of a user interacting with an LLM chatbot for health to understand conditions for acceptance, useability, and believability. We provided versions of these scenarios that were augmented to include location, gender and religion, to understand whether the above metrics changed with additional contextual responses. Finally we asked respondents to indicate specific non-personal hypothetical questions they would ask an AI chatbot about health. The full set of questions asked in the survey can be found in Appendix A.3.1.
3.1.3. Statistical analysis
@@
3.2. Expert in-depth interviews
3.2.1. IDI Expert Participant Selection
@@
To recruit experts, we reached out to people in our networks using an email draft which included a link to fill out a participation interest survey for the interviews. We also shared the study description and fliers with a link/QR code to the form in AI communities and group chats, and on social media pages including LinkedIn. Participants were recruited between October 24 2023 and January 30th 2024.
@@
Since the in-depth interviews (IDIs) specifically focused on experts, we recruited participants who had expertise in AI and health, practicing health providers, public health researchers, health policy experts, and entrepreneurs. We had 445 interested participants sign up to take part in the study. We went through five rounds of short-listing to narrow down the participants based on responses in the participant interest survey. Participants were selected based on (1) years of experience, prioritizing >10 years of experience, followed by 5+ years of experience in their respective fields, (2) field of practice–we ensured there was a balanced representation of clinicians, public health, and AI researchers, (3) focus of research/work on Africa, and (4) country of residence, prioritizing African residents. We also included non-African residents who were African diaspora and/or had >10 years of experience working in Africa. Demographics of the participants as well as relevant criteria such as years of experience can be found in Table 3. Expertise in AI Trust/safety/responsible AI was not a selection criteria for experts, and none of the experts we interviewed indicated specifically that they were in the field of AI Trust/safety/responsible AI.
3.2.2. IDI Procedure
@@
After agreeing to informed consent, the participants filled out a brief demographic survey. The interview was then scheduled and then conducted online over Google Video Conference (GVC) and audio-recorded. There were two individual moderators throughout the total qualitative interviews, and each interview was conducted with 1 moderator and 1 participant. During the interview, a moderator engaged with the participants in in-depth, semi-structured interviews that lasted approximately 90-120 minutes. Before each interview, the interviewer thoroughly explained the interview process, emphasized confidentiality and addressed any questions or concerns raised by the participants. After each interview, the voice recordings were saved as MP3 files and no identifying data (i.e., participants’ imagery) was retained. Voice recordings were transcribed by a vendor (D’ Well), and the transcripts stored in a secure data bucket without any PII. The full interview guide can be found in Appendix A.3.2.
@@
The focus of the interviews explored the participants’ perspectives on (a) equity in healthcare; (b) historical biases in health; (c) AI in healthcare; (d) algorithmic fairness framing and applications; (e) stereotypes, sensitive attributes, and vulnerable sub-populations; and (f) an understanding of existing definitions of fairness. Further, we explored (g) perceptions of colonialism and the Global North/South divide.
3.2.3. Data Analysis